6x-4x+x=18 what does x equal show work thanks
Answer by funmath(2933)

6x-4x+x=18
(6-4+1)x=18
3x=18
3x/3=18/3
x=6
Check by substituting the answer you got in for x:
6(6)-4(6)+6=18
36-24+6=18
12+6=18
18=18 We're right!!!!!
x=6